OpenStudy (anonymous):

\[6t ^{-2}-29t ^{-1}+20=0\]

OpenStudy (anonymous):

gimmick is to solve this as a quadratic equation is \[t^{-1}\]

OpenStudy (anonymous):

in other words solve \[6x^2-29x+20=0\] \[(x-4)(6x-5)=0\] \[x=4,x=\frac{5}{6}\] \[t^{-1}=4,t=\frac{1}{4}\] \[t^{-1}=\frac{6}{5}\] \[t=\frac{5}{6}\]

OpenStudy (anonymous):

Thank you very much..